OpenStudy (anonymous):

Another one: Given that all roots of the equation \( x^5 − 10x^4 + ax^3 + bx^2 + cx − 32 = 0 \) are real and positive.Find \( a+b+c \)

OpenStudy (anonymous):

40 LOL

OpenStudy (anonymous):

Explain.

OpenStudy (anonymous):

from viete i got that x1*x2*x3*x4*x5=-32 so i just guess that 2 is one of the factor and i guess by (x-2)^5 and multiply it together and get the result \[x^5-10x^4+40x^3-80x^2+80x-32=0\] so a+b+c=40-80+80=40

OpenStudy (anonymous):

could be a kind of stupid way to do lol

OpenStudy (anonymous):

lol, yeah but you got the correct solution.

OpenStudy (anonymous):

cool lol you have a better way?

OpenStudy (anonymous):

May be ;), but let see how others solve it

OpenStudy (anonymous):

wow thats totally lucky.